solve for x: 0.008 -0.02x=0.03 Then evaluate : x^2+x+.89
Answer by Boreal(15235)

0.008 -0.02x=0.03 Then evaluate : x^2+x+.89
-0.02x=0.024
divide both sides by -0.02
x= -1.2
======
x^2+x+89, where x=-1.2
1.44-1.2+0.89=2.33-1.2=1.13